owl shift summary 6/8/99

S. Elizabeth Turner (eturner@kudzu)
Tue, 8 Jun 1999 07:06:19 -0500

On Shift: Elizabeth, Jason L.
Others Present: Rick K, Val, Ed

Hours of beam delivered at all intensities: ~7
Hours of beam delivered at nominal intensity(3-4 E11): ~7
Hours of E832 data (special runs not included): ~2.5
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time: front end trouble, loading new CsI HVs

========================================================================

Runs ended this shift:

13425: BEAM_832_ECAL ~2 million electrons
13426: BEAM_832 (stopped due to front end trouble)
13430: BEAM_832 (stopped to modify dest. logic)
13431: BEAM_832 (stopped to change CsI HVs)
13433: BEAM_832_ECAL in progress

Tapes brought to FCC this shift: 0

========================================================================

Accesses during shift:

Rick K. and Jason L. make access attempting to debug front end.

========================================================================
Description of Shift:

Plans for this shift were to finish collecting 2 million ECAL
triggers for CsI, to do 832 trigger studies, then to take 1 million
more ECAL triggers after changing HVs.

DETAILS of SHIFT:
------------------
00:30 Rick changes REG readout threshold from <PED>+4 counts to <PED>+10 counts.
Notes that offline cut is 2 MIPs (~60 counts) which is far from threshold.

00:51 Lost 10 spills (beginning run 13425 spill 150) due to FERA peds busy.

1:15 Stop ECAL run 13425 with 2 million ECAL tags.

1:15 BEAM_832 trigger studies (results in KTeV99 E832 rates notebook)

2:40 Start run 13426 (BEAM_832). Val notes that runs 13245 & 13246 should
not be used for offline analysis because Q/E constants were back-dated
after offline calibration.

2:50 Jason L. swaps some incorrect banana cabling.

3:10 Front end troubles begin. Symptoms were missing FFFF words and unknown
crates. Tried multiple kfrends, cycled power on repeater, Jason and Rick
went on access while running random pulsers. After cycling power on repeater
again, everything is fine. Note that sometimes repeater power must be
cycled more than once to fix problems.

5:00 Stopped run 13430 to modify destination logic. Init run 13431, writing RANL3
to disk to study tape balancing.

5:00 Find that KQ/BAN L2 eff = 76% for BEAM_832 run 13426,13430, 13431.

5:25 Stop run 13431 to reload HV for CsI and REG after calibration and to
study KQ/BAN L@ problem.

5:30 Discover that KQ/BAN efficiency problem is due to bad AZLAT, trigger is ok.

6:00 Load REG and CsI HVs before start of run 13432.

6:50 Start run 13433

6:55 ET problem, any data from first 5 spills of run 13433 is no good.